OpenStudy (elleblythe):

How to solve for: 16^x-4^x=6

OpenStudy (anonymous):

take the ln for the eq.

OpenStudy (anonymous):

xln(16)-xln(4)=ln(6) x(ln(16/4))=ln(6)

OpenStudy (dan815):



OpenStudy (owlfred):

Hmm, no

OpenStudy (owlfred):

\[\ln(a - b) \ne \ln a - \ln b\]

OpenStudy (dan815):